OpenStudy (anonymous):

Please help? Find sin θ if cot θ = - 2 and cos θ < 0.

OpenStudy (goformit100):

Take trigonometry in use

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

cos theta<0 means cos theta is negative so theta lies in 2nd or 3rd quadrANT. cot theta is negative so theta lies in 2nd or 4th quadrant. we conclude theta lies in 2nd quadrant. csc^2 theta-cot^2 theta=1 csc^2 theta=1+cot^2 theta=1+4=5 sin^2 theta=1/5 sin theta=1/sqrt5 (in 2nd quadrant sin theta is positive)

OpenStudy (anonymous):

My answer choices were: -1/2, -5, sqrt5/2, and sqrt5/5..

OpenStudy (anonymous):

@plohrr

OpenStudy (anonymous):

yes surgithayer was right. i don't understand why its not a choice.

OpenStudy (anonymous):

maybe we need to take the radical out? in that case it would be sqrt5/2

OpenStudy (anonymous):

Oh, okay. That makes more sense! Thank you @plohrr ! :D God bless you!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Nope, that wasn't right.. :( I missed that one, and another one.

OpenStudy (anonymous):

… im afraid i can't help with this one anymore. sorry :(

OpenStudy (anonymous):

Crap..

OpenStudy (anonymous):

if sqrt5/2 isn't right, i don't know what is lol

OpenStudy (anonymous):

@jim_thompson5910 Please help me! :c

jimthompson5910 (jim_thompson5910):

See attached (for some reason the formulas aren't displaying properly)

OpenStudy (anonymous):

Oh, thank you so much! Wonderful. :)